# KDP as a thermal blocking filter Deep near IR observations with a warm narrow band filter

### v1

@@ Abstract
Ground-based astronomy suffers from strong atmospheric line- and thermal continuum emission, at the near infrared (NIR, 0.7–1.1 ), and short-wave infrared (SWIR, 1.1–2.5 ) wavelengths. The thermal continuum emission increases exponentially towards the red sensitivity cutoff of the state-of-the-art 2.5 cutoff SWIR detectors. Given availability of an optical quality shortpass filter material with strong blocking density in the SWIR, lower cost instrumentation, and higher performance filters could be designed. We demonstrate monopotassium dihydrogen phosphate (KDP, chemical formula ) as a strong candidate for this purpose. KDP is fully transparent at wavelengths from ultraviolet to 1.3 , but becomes highly opaque at wavelengths 2 . We demonstrate on-sky use of KDP by improving performance of a cryogenic broadband filter with known off-band thermal leak, and using a non-cryogenic narrow band filter for deep observation. KDP reduces the sky background by 4.5 magnitudes in the leaky Z-band filter we use. Our 4 nm wide, central wavelength 1.191 narrowband filter in combination with KDP reduces the sky surface brightness by three magnitudes compared to a J broadband. We find a sky surface brightness of 18.5 mag arcsec-2 at 1.191 , and interpret it as the airglow continuum. KDP is an excellent thermal blocker, when its temperature is maintained above its Curie point at 123 K. Below Curie point, KDP transforms its crystal structure, degrading its otherwise good imaging properties.
@@ 1 Introduction
The use of Potassium Dihydrogen Phosphate (KDP, ) for blocking thermal radiation has been suggested in the literature: to let the 2.5 cutoff detectors be used without a cryogenic optical train (P. J. Amado et al., 2012). We have not seen a study applying KDP as a thermal radiation blocking filter in the context of astronomical observations. In this work, we will do exactly that and report on our findings. KDP and its deutered variant Potassium Dideuterium Phosphate (D-KDP, ) are commonly used non-linear optical crystals in laser frequency modulation. Historically, KDP and D-KDP have been studied extensively in the context of high-power lasers for fusion power (D. Eimerl, 1987; J. J. De Yoreo et al., 2002). In addition to the D-KDP, KDP has a few other isomorphic crystals, which share similar physical properties.
One of the unique physical properties of KDP is its high absorbance in the SWIR wavelength regime, and very high transmission from UV to 1.3 (Fig. 1), making it an excellent thermal radiation blocking filter. This is notable since only a few optical materials exist that absorb in the thermal infrared, while transmitting in VIR–NIR range (D. Eimerl, 1987; M. J. Weber, 2020). Available thermal radiation absorbing glasses do not offer nearly as high optical density (OD) in the SWIR and they also absorb at shorter wavelengths, making them less ideal as short-pass filters. KDP’s isomorphic crystals (Table 1) show similar properties of being practically fully transparent at wavelengths shorter than their shortpass cutoff wavelength, and very high optical density at wavelengths longer than their cutoff wavelength. Several immediately interesting use cases can be realized, including:
@@ 2 Experimental setup
We used the Nordic Optical Telescope’s (NOT) NOTCam instrument (T. M. Abbott et al., 2000) as a test bed. NOTCam is an imager-spectrograph designed to operate in the SWIR wavelength region including the K band. It has a hemispherical entrance baffle, and a cold Lyot stop limiting off-axis thermal background. The instrument temperature is at 73 K. NOTCam has an aperture wheel, two filter wheels, a stop wheel, and a grism wheel (Fig. 5). We use the filter wheels for mounting the broadband filters, and the aperture wheel for mounting the KDP window. Additionally, the warm narrowband filter was installed in front of NOTCam’s cryostat window in the telescope converging beam, i.e. in fully ambient, non-cryogenic conditions.
2.1 KDP
@@
KDP and its deutered variant, D-KDP, are commonly used for frequency modulation in NIR lasers. Possibly, most well known, KDP and D-KDP have been used in the high power lasers of the National Ignition Facility at Lawrence Livermore National Laboratory (J. J. De Yoreo et al., 2002). KDP can be grown to large size, growing it is relatively quick, and it can be polished to optical quality, making it a low cost and easy to work with optical material. KDP crystals can be doped with different impurities, and research on KDP continues today in the field of solid state dye lasers.
@@
Isomorphic crystals of KDP include Ammonium- (ADP), Rubidium- (RDP) and Cesium- (CDP) Dihydrogen Phosphates, and similarly, Potassium (KDA), Arsenic- (ADA), Rubidium (RDA) and Cesium (CDA) arsenates, and their deutered counterparts (e.g. Ammonium Dideuterium Phosphate, D-ADP). KDP and its isomorphisms are polymorphic crystals which exhibit interesting dielectric and non-linear optical properties. We will focus on two of their physical properties: the spectral transmission, and the Curie temperature .
@@
KDP and its isomorphisms are nearly completely opaque in the thermal infrared, while being fully transparent at wavelengths shorter than 1.3 (D. Eimerl, 1987), making them ideal short-pass ’filter glasses’, thermal blockers, for NIR filter design. Being polymorphic crystals, KDP and its isomorphisms change from paramagnetic to ferromagnetic state when cooled below their . Cooling them below also changes the crystal structure, which in turn introduces unwanted optical qualities: at temperatures below the , we see position-dependent transmission non-uniformity and possible polarization-induced point spread function distortions with our z-cut KDPs (Fig. 2). We have observed similar phase-transition effects in both KDP and D-DKP.
@@
Because of the phase transition, KDP and its isomorphs have to be operated above their to maintain the desired optical qualities. In return, this poses a trade-off between the external thermal flux absorbed by the crystal, and the thermal flux emitted by the crystal itself. The is a key property to balance the trade-off. This generally favors materials with lower for SWIR astronomy applications, suggesting the most promising materials to be KDP, KDA and RDA. We summarize the KDP isomorphisms and their Curie temperatures, and approximate short-pass cutoff wavelengths () in Table 1. We refer to the optical and mechanical properties of KDP reported by C. C. Stephenson & J. G. Hooley (1944); F. Zernike (1964); D. Eimerl (1987); W. M. Haynes et al. (2016); M. J. Weber (2020).
@@
We purchased three z-cut KDP and one D-KDP crystal from United Crystals Inc. for this study. Two Ø25 mm diameter and 4 mm thick KDPs, and one Ø25 mm diameter and 4 mm thick D-KDP were already purchased in 2016. The first observations followed soon after. The phase transition at was not immediately understood, and the project fell dormant for a while for that reason. When it became clear that the change in optical properties was due to the phase transition, calculations were done on the reachable equilibrium temperature inside NOTCam (Sec. 4.2). A third Ø10 mm diameter and 0.5 mm thick KDP window was purchased and used for the observations reported in this work.
@@
The KDPs and the D-KDP were installed in NOTCam’s aperture wheel (Fig. 5). The 4 thick KDPs and the D-KDP were installed in thermal contact with the instrument, while 0.5 thick KDP was prepared with a thermally insulating glass fiber holder. The KDPs in the aperture wheel are placed in the focal plane of the telescope, allowing us to warm the thermally insulated KDP above the by the thermal radiation from the vacuum window (Sec. 4.2).
@@
Note. — Curie temperatures and cutoff wavelengths for ordinary , and extraordinary beam have been collected from D. Eimerl (1987), and are measured for 11 mm thick samples. We define the as a 50% falling edge on the transmission curve. D-* indicates the deutered variants of Potassium (K), Arsenic (A), Rubidium (R), and Cesium (C) phosphates (P), and arsenates (A).
2.2 The warm narrow-band filter
@@
Manufacturing high performance interference filters is challenging and expensive, and a trade-offs between peak efficiency, in-band transmission, and off-band blocking must be made. In the SWIR, a large spectral range is covered, leaving observations vulnerable to off-band, usually thermal, leaks. The off-band contribution can be significant, especially in narrow-band photometry, where the flux in the band-pass is weak, degrading the photometric performance. Due to the ability of KDP to absorb thermal radiation, filters with band-passes shorter than 1.3 may be placed in ambient (warm) conditions. A KDP thickness of 1 mm is sufficient to absorb thermal emission from the ’warm’ filter, to a level that has no impact on the observations.
@@
We designed a custom warm NB filter, to demonstrate this. The filter is designed to be used in combination with the NOTCam the J-band filter and a 0.5 mm KDP. The NB filter is only blocking off-band light through the J-band (Fig. 3). The NB filter has a central wavelength of 1.191 , and full-width-at-half-maximum of 4 nm (R300), being located in the largest gap between atmospheric hydroxyl (OH) emission lines within the J band. The filter was designed and manufactured by Ferroperm Optics A/S. We will refer it as BP1191-4.
@@
The BP1191-4 filter is based on a Fabry-Pérot type design with four cavities having a theoretical bandwidth of 4.2 nm on the front surface of a Borofloat 33-substrate. A stress compensating anti-reflection coating was added on the rear surface to increase the transmission. The filter was manufactured in a Generation III Helios 800-chamber, employing the Plasma Assisted Reactive Magnetron Sputtering (PARMS) coating technology. Niobium pentoxide, Nb2O5, was used as a high refractive index and silicon dioxide, SiO2, as a low refractive index material. The monitoring strategy of the bandpass filter was direct monitoring at the center wavelength on a Borofloat 33-substrate utilizing the self-compensating turning point method (P. W. Baumeister, 2004). The result of the production coating runs was a 1191 nm filter with a bandwidth of 3.9 nm and a peak transmission of 94% measured on a Perkin Elmer L1050 -spectrophotometer (Fig. 3).
2.3 Broad-band filters
@@
The NOTCam broadband J, H, K filters were ordered as a part of the Mauna Kea Observatories near-infrared (MKO-NIR) set and have well defined bandpasses (D. A. Simons & A. Tokunaga, 2002; A. T. Tokunaga et al., 2002; A. T. Tokunaga & W. D. Vacca, 2005). Two copies of the filters exist at the NOT: one installed inside NOTCam, and another stored in a clean-room outside of the instrument. For this study, we scanned the spare J-band filter, using a Perkin-Elmer L1050 -spectrophotometer at Ferroperm Optics A/S in Vedbæk, Denmark. The spectral range 0.3 to 2.7 was covered, with the motivation of understanding the off-band blocking characteristics of the filter.
@@
NOTCam’s Z- and Y-band filters were intended to have the corresponding UKIDSS bandpasses (P. C. Hewett et al., 2006). The filters were not manufactured accordingly however, and especially the Z-band filter has pour off-band blocking. The Z and Y band filters were scanned at the Instituto Astrofisica de Canarias (IAC) in range 0.3 to 2.6 . The Z-band filter was confirmed to have a significant thermal leak (Fig. 4), making it an ideal optical component for demonstrating the performance of KDP as a thermal blocker.
2.4 Detector
@@
NOTCam is equipped with a 1024 1024 18.5 pixel HgCdTe/Al2O3 HAWAII PACE I -detector with a cutoff at 2.5 (L. J. Kozlowski et al., 1994), which is controlled by a Brorfelde Image Acquisition System (BIAS) -controller (J. Klougart, 1995). With the NOTCam wide-field camera, the detector covers a 4′ 4′ area on sky. The BIAS controller offers two read modes, 1.) reset-read-read (correlated double sampling), and 2.) ramp-sample (non-destructive readout with specified time interval). Due to controller-related specifics not known to us, images from the NOTCam detector show more complex DC structure than other similar detectors. The phenomena is related to the first few reads, and in case of ramp-sampling data can be circumvented by re-fitting the ramps in postprocessing: instead of using the controller provided ramp-sample fits, we have re-fit all data used in this work, ignoring the first read in a sequence.
@@
Unlike typical SWIR observations, we are detector noise limited in the shorter integrations in the BP1191-4 bandpass. With the BP1191-4 filter, we readout the detector in ramp-sampling mode with 14 samples per integration, and in all broadbands, with as many 4 s sub-reads as it is possible while keeping the total integration within the 1% linearity limit of the detector. We do not have a measure of the inter-pixel capacitance (A. C. Moore et al., 2006), and to avoid uncertainty in the gain conversion by determining, and applying, all zeropoints in terms of Analog-to-Digital-Units (ADUs). NOTCam gain measured with the photon transfer tech (J. R. Janesick et al., 1987) is 2.6 e- ADU-1. We use this gain value when estimating the thermal flux from KDP.
@@ 3 Observations
A second observing campaign was started in 2025, after the thermally insulated 0.5 mm KDP window was installed in NOTCam. Sky conditions on 2025 January 14 and 15 were excellent. New observations were done when the BP1191-4 filter became available in 2025 June. The BP1191-4 bandpass is bue enough that scattered Moon light would affect observations, thus they have to be obtained under dark sky conditions. The observations on 2025 June 30 were made under excellent conditions. The fields observed in BP1191-4 were re-observed in J band on 2025 August 8, but these observations suffered from large amounts of Saharan dust, so called ”Calima”. Since the 2025 June 30 observations were detector systematics limited (Sec. 2.4), one more set of BP1191-4 observations was made on 2025 August 23, unfortunately also suffering from Saharan dust.
@@
References. — Gemini UKIRT MKO photometric standard stars catalogue, S. K. Leggett et al. (2006).
@@
4 Analysis and results
4.1 KDP transmission and optical density

### v2

@@ Abstract
Ground-based astronomy suffers from strong atmospheric line- and thermal continuum emission, at the near infrared (NIR, 0.7–1.1 ), and short-wave infrared (SWIR, 1.1–2.5 ) wavelengths. The thermal continuum emission increases exponentially towards the red sensitivity cutoff of the state-of-the-art 2.5 cutoff SWIR detectors. Given availability of an optical quality shortpass filter material with strong blocking density in the SWIR, lower cost instrumentation, and higher performance filters could be designed. We demonstrate monopotassium dihydrogen phosphate (KDP, chemical formula ) as a strong candidate for this purpose. KDP is fully transparent at wavelengths from ultraviolet to 1.3 , but becomes highly opaque at wavelengths 2 . We demonstrate on-sky use of KDP by improving performance of a cryogenic broadband filter with known off-band thermal leak, and using a non-cryogenic narrow band filter for deep observation. KDP reduces the sky background by 4.5 magnitudes in the leaky Z-band filter we use. Our 4 nm wide, central wavelength 1.191 narrowband filter in combination with KDP reduces the sky surface brightness by three magnitudes compared to a J broadband, although the effect of KDP is minor due to high blocking density of the broaband filter. We find a sky surface brightness of 18.5 mag arcsec-2 in our bandpass at 1.191 . KDP is an excellent thermal blocker, when its temperature is maintained above its Curie point at 123 K. Below Curie point, KDP transforms its crystal structure, degrading its otherwise good imaging properties.
@@ I Introduction
The use of Potassium Dihydrogen Phosphate (KDP, ) for blocking thermal radiation has been suggested in the literature: to let the 2.5 cutoff detectors be used without a cryogenic optical train (Amado12). We have not seen a study applying KDP as a thermal radiation blocking filter in the context of astronomical observations. In this work, we will do exactly that and report on our findings. KDP and its deuterated variant Potassium Dideuterium Phosphate (D-KDP, ) are commonly used non-linear optical crystals in laser frequency modulation. Historically, KDP and D-KDP have been studied extensively in the context of high-power lasers for fusion power (Eimerl87; Yoreo02). In addition to the D-KDP, KDP has a few other isomorphic crystals, which share similar physical properties.
One of the unique physical properties of KDP is its high absorbance in the SWIR wavelength regime, and very high transmission from UV to 1.3 (Fig. 1), making it an excellent thermal radiation blocking filter. This is notable since only a few optical materials exist that absorb in the thermal infrared, while transmitting in VIR–NIR range (Eimerl87; CRClaser). Available thermal radiation absorbing glasses do not offer nearly as high optical density (OD) in the SWIR and they also absorb at shorter wavelengths, making them less ideal as short-pass filters. KDP’s isomorphic crystals (Table 1) show similar properties of being practically fully transparent at wavelengths shorter than their shortpass cutoff wavelength, and very high optical density at wavelengths longer than their cutoff wavelength. Several immediately interesting use cases can be realized, including:
@@ II Experimental setup
We used the Nordic Optical Telescope’s (NOT) NOTCam instrument (Abbot00) as a test bed. NOTCam is an imager-spectrograph designed to operate in the SWIR wavelength region including the K band. It has a hemispherical entrance baffle, and a cold Lyot stop limiting off-axis thermal background. The instrument temperature is at 73 K. NOTCam has an aperture wheel, two filter wheels, a stop wheel, and a grism wheel (Fig. 5). We use the filter wheels for mounting the broadband filters, and the aperture wheel for mounting the KDP window. Additionally, the warm narrowband filter was installed in front of NOTCam’s cryostat window in the telescope converging beam, i.e. in fully ambient, non-cryogenic conditions.
II.1 KDP
@@
KDP and its deutered variant, D-KDP, are commonly used for frequency modulation in NIR lasers. Possibly, most well known, KDP and D-KDP have been used in the high power lasers of the National Ignition Facility at Lawrence Livermore National Laboratory (Yoreo02). KDP can be grown to large size, growing it is relatively quick, and it can be polished to optical quality, making it a low cost and easy to work with optical material. KDP crystals can be doped with different impurities, and research on KDP continues today in the field of solid state dye lasers.
@@
Isomorphic crystals of KDP include Ammonium- (ADP), Rubidium- (RDP) and Cesium- (CDP) Dihydrogen Phosphates, and similarly, Potassium (KDA), Arsenic- (ADA), Rubidium (RDA) and Cesium (CDA) arsenates, and their deuterated counterparts (e.g. Ammonium Dideuterium Phosphate, D-ADP). KDP and its isomorphisms are polymorphic crystals which exhibit interesting dielectric and non-linear optical properties. We will focus on two of their physical properties: the spectral transmission, and the Curie temperature .
@@
KDP and its isomorphisms are nearly completely opaque in the thermal infrared, while being fully transparent at wavelengths shorter than 1.3 (Eimerl87), making them ideal short-pass ’filter glasses’, thermal blockers, for NIR filter design. Being polymorphic crystals, KDP and its isomorphisms change from paramagnetic to ferromagnetic state when cooled below their . Cooling them below also changes the crystal structure, which in turn introduces unwanted optical qualities: at temperatures below the , we see position-dependent transmission non-uniformity and possible polarization-induced point spread function distortions with our z-cut KDPs (Fig. 2). We have observed similar phase-transition effects in both KDP and D-DKP.
@@
Because of the phase transition, KDP and its isomorphs have to be operated above their to maintain the desired optical qualities. In return, this poses a trade-off between the external thermal flux absorbed by the crystal, and the thermal flux emitted by the crystal itself. The is a key property to balance the trade-off. This generally favors materials with lower for SWIR astronomy applications, suggesting the most promising materials to be KDP, KDA and RDA. We summarize the KDP isomorphisms and their Curie temperatures, and approximate short-pass cutoff wavelengths () in Table 1. We refer to the optical and mechanical properties of KDP reported by Stephenson44a; Zernike64; Eimerl87; CRCchem; CRClaser.
@@
We purchased three uncoated z-cut KDP and one D-KDP crystals from United Crystals Inc. for this study. Two Ø25 mm diameter and 4 mm thick KDPs, and one Ø25 mm diameter and 4 mm thick D-KDP were already purchased in 2016. The first observations followed soon after. The phase transition at was not immediately understood, and the project fell dormant for a while for that reason. When it became clear that the change in optical properties was due to the phase transition, calculations were done on the reachable equilibrium temperature inside NOTCam (Sec. IV.2). A third Ø10 mm diameter and 0.5 mm thick KDP window was purchased and used for the observations reported in this work.
@@
The KDPs and the D-KDP were installed in NOTCam’s aperture wheel (Fig. 5). The 4 thick KDPs and the D-KDP were installed in thermal contact with the instrument, while 0.5 thick KDP was prepared with a thermally insulating glass fiber holder. The KDPs in the aperture wheel are placed in the focal plane of the telescope, allowing us to warm the thermally insulated KDP above the by the thermal radiation from the vacuum window (Sec. IV.2).
@@
Note. — Curie temperatures and cutoff wavelengths for ordinary , and extraordinary beam have been collected from Eimerl87, and are measured for 11 mm thick samples. We define the as a 50% falling edge on the transmission curve. D-* indicates the deuterated variants of Potassium (K), Arsenic (A), Rubidium (R), and Cesium (C) phosphates (P), and arsenates (A).
II.2 The warm narrow-band filter
@@
Manufacturing high performance interference filters is challenging and expensive, and a trade-off between peak efficiency, in-band transmission, and off-band blocking must be made. In the SWIR, a large spectral range is covered, leaving observations vulnerable to off-band, usually thermal, leaks. The off-band contribution can be significant, especially in narrow-band photometry, where the flux in the band-pass is weak, degrading the photometric performance. Due to the ability of KDP to absorb thermal radiation, filters with band-passes shorter than 1.3 may be placed in ambient (warm) conditions. A KDP thickness of 1 mm is sufficient to absorb thermal emission from the warm filter, to a level that has no impact on the observations.
@@
We designed a custom warm NB filter to be used in combination with the NOTCam J-band filter and a 0.5 mm KDP. The KDP is to ensure that there will be no thermal leak through the broadband filter. The NB filter is only blocking off-band light through the J-band (Fig. 3), and it has a central wavelength of 1.191 , with full-width-at-half-maximum of 4 nm (R300). The bandpass is located in the largest gap between atmospheric hydroxyl (OH) emission lines within the J band. The filter was designed and manufactured by Ferroperm Optics A/S. We will refer it as BP1191-4.
@@
The BP1191-4 filter is based on a Fabry-Pérot type design with four cavities having a theoretical bandwidth of 4.2 nm on the front surface of a Borofloat 33-substrate. A stress compensating anti-reflection coating was added on the rear surface to increase the transmission. The filter was manufactured in a Generation III Helios 800-chamber, employing the Plasma Assisted Reactive Magnetron Sputtering (PARMS) coating technology. Niobium pentoxide, Nb2O5, was used as a high refractive index and silicon dioxide, SiO2, as a low refractive index material. The monitoring strategy of the bandpass filter was direct monitoring at the center wavelength on a Borofloat 33-substrate utilizing the self-compensating turning point method (Baumeister04). The result of the production coating runs was a 1191 nm filter with a bandwidth of 3.9 nm and a peak transmission of 94% measured on a Perkin Elmer L1050 -spectrophotometer (Fig. 3).
II.3 Broad-band filters
@@
The NOTCam broadband J, H, K filters were ordered as a part of the Mauna Kea Observatories near-infrared (MKO-NIR) set and have well defined bandpasses (Simons02; Tokunaga02; Tokunaga05). Two copies of the filters exist at the NOT: one installed inside NOTCam, and another stored in a clean-room outside of the instrument. For this study, we scanned the spare J-band filter, using a Perkin-Elmer L1050 -spectrophotometer at Ferroperm Optics A/S in Vedbæk, Denmark. The spectral range 0.3 to 2.7 was covered, with the motivation of understanding the off-band blocking characteristics of the filter.
@@
NOTCam’s Z- and Y-band filters were intended to have the corresponding UKIDSS bandpasses (Hewett06). The filters were not manufactured accordingly however, and especially the Z-band filter has poor off-band blocking. The Z and Y band filters were scanned at the Instituto Astrofisica de Canarias (IAC) in range 0.3 to 2.6 . The Z-band filter was confirmed to have a significant thermal leak (Fig. 4), making it an ideal optical component for demonstrating the performance of KDP as a thermal blocker.
II.4 Detector
@@
NOTCam is equipped with a 1024 1024 18.5 pixel HgCdTe/Al2O3 HAWAII PACE I -detector with a cutoff at 2.5 (Kozlowski94), which is controlled by a Brorfelde Image Acquisition System (BIAS) -controller (Klougart95). With the NOTCam wide-field camera, the detector covers a 4′ 4′ area on sky. The BIAS controller offers two read modes, 1.) reset-read-read (correlated double sampling), and 2.) ramp-sample (non-destructive readout with specified time interval). Due to controller-related specifics not known to us, images from the NOTCam detector show more complex DC structure than other similar detectors. The phenomena is related to the first few reads, and in case of ramp-sampling data can be circumvented by re-fitting the ramps in postprocessing: instead of using the controller provided ramp-sample fits, we have re-fit all data used in this work, ignoring the first read in a sequence.
@@
Unlike typical SWIR observations, we are detector noise limited in the shorter integrations in the BP1191-4 bandpass. With the BP1191-4 filter, we readout the detector in ramp-sampling mode with 14 samples per integration, and in all broadbands, with as many 4 s sub-reads as it is possible while keeping the total integration within the 1% linearity limit of the detector. We do not have a measure of the inter-pixel capacitance (Moore06), and to avoid uncertainty in the gain conversion by determining, and applying, all zeropoints in terms of Analog-to-Digital-Units (ADUs). NOTCam gain measured with the photon transfer tech (Janesick87) is 2.6 e- ADU-1. We use this gain value when estimating the thermal flux from KDP.
@@ III Observations
A second observing campaign was started in 2025, after the thermally insulated 0.5 mm KDP window was installed in NOTCam. Sky conditions on 2025 January 14 and 15 were excellent. New observations were done when the BP1191-4 filter became available in 2025 June. The BP1191-4 bandpass is blue enough that scattered Moon light would affect observations, thus they have to be obtained under dark sky conditions. The observations on 2025 June 30 were made under excellent conditions. The fields observed in BP1191-4 were re-observed in J band on 2025 August 8, but these observations suffered from large amounts of Saharan dust, so called ’Calima’. Since the 2025 June 30 observations were detector systematics limited (Sec. II.4), one more set of BP1191-4 observations was made on 2025 August 23, unfortunately also suffering from Saharan dust.
@@
References. — Gemini UKIRT MKO photometric standard stars catalogue, Legget06.
@@
IV Analysis and results
IV.1 KDP transmission and optical density